# Explaining Tournament Solutions with Minimal Supports

### v2

@@ 11          $\mu_{\mathcal{X}}(c^{\prime},c)\leftarrow n-t$
Let $\mathcal{X}$ be a SMS for $w\in\operatorname{BO}(G)$. With Lemma 27, we show that if $\sigma_{w}\leq n(m-1)-\min_{c\neq w}\mu(w,c)$ then the lower bound of Lemma 24 is reached and for all SMSs $\mathcal{X}$ for $w\in\operatorname{BO}(G)$, $|\mathcal{X}|{=}(m-1)(n(m-1)-\sigma_{w})$.
@@
The lower bound directly result from the Proposition 15 as $n(m-1)-\lfloor n(1-1/m)\rfloor=n(m-1)-\left\lfloor n(m-1)/m\right\rfloor=\left\lceil n(m-1)-n(m-1)/m\right\rceil=\left\lceil\frac{n(m-1)^{2}}{m}\right\rceil$. For the upper bound, note that the Borda score of the winner $\sigma_{w}$ is greater than the average score of all candidates, i.e., $\sigma_{w}\geq\frac{1}{m}n\frac{m(m-1)}{2}$. In the worst case $\sigma_{w}=\left\lceil n\frac{m-1}{2}\right\rceil$ and $|\mathcal{X}|=(m-1)(n(m-1)-\left\lceil n\frac{m-1}{2}\right\rceil)=(m-1)\left\lfloor n\frac{m-1}{2}\right\rfloor$. ∎
@@
To compute a maxwin-SMS in $\mathcal{O}(m^{2}\log n)$, we propose the following Algorithm 2 which can naturally be derived from the proof of Porposition 15 on the size of SMSs. ∎
@@ 20    $\text{rmed}\leftarrow\text{rmed}+\text{iteRm}$
Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},E)$ , and a winning candidate $w\in\operatorname{BO}(G)$, for all SMS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$,
@@
    else if $\exists!c_{0}\in\mathcal{C}\setminus{\{w\}}$ s.t. $\delta(c_{0},{E}^{w\rightarrow})\leq 0$ then $|{E}^{w\rightarrow}\setminus{\mathcal{X}}|\leq\min_{c\in\mathcal{C}\setminus{\{w,c_{0}\}}}\delta(c,{E}^{w\rightarrow})$ and exists a SMS containing ${E}^{w\rightarrow}$
@@
    else for all SMS $\mathcal{Y}$ for $w\in\operatorname{BO}(G)$, $|\operatorname{WC}(\mathcal{X})-\operatorname{WC}(\mathcal{Y})|\leq 1$.
@@
(i) We proceed by contradiction. Suppose, there is a SMS $\mathcal{X}$ of $w\in\operatorname{BO}(G)$ such that ${E}^{w\rightarrow}\not\subseteq\mathcal{X}$. Then, $\exists c_{2}\in\mathcal{C}\setminus{\{w\}}$ such that $(w,c_{2})\in E$ and $(w,c_{2})\not\in\mathcal{X}$. Additionally, we know that $\exists e_{0}\in{\mathcal{X}}^{\rightarrow c_{0}}$ such that $e_{0}\neq(w,c_{0})$ (all $c_{0}$’s defeats in $\mathcal{X}$ cannot come from $w$ since $\delta(c_{0},{E}^{w\rightarrow})<0$). Similarly, $\exists e_{1}\in{\mathcal{X}}^{\rightarrow c_{1}}$, $e_{1}\neq(w,c_{1})$. Let $\mathcal{X}^{\prime}=\mathcal{X}\setminus{\{e_{0},e_{1}\}}\cup\{(w,c_{2})\}$. $\forall c\in\mathcal{C}\setminus{\{w,c_{0},c_{1},c_{2}\}},\,\delta(c,\mathcal{X}^{\prime})=\delta(c,\mathcal{X})+1\geq 0$. $\delta(c_{2},\mathcal{X}^{\prime})=\delta(c_{2},\mathcal{X})+1\geq 0$. $\forall c\in\{c_{0},c_{1}\}\setminus{\{c_{2}\}}$, $\delta(c,\mathcal{X}^{\prime})=\delta(c,\mathcal{X})-1+1\geq 0$. Hence $\mathcal{X}^{\prime}$ is a weak MS and it contains at least one MS with a smaller cardinality. However, $|\mathcal{X}^{\prime}|<|\mathcal{X}|$ and $\mathcal{X}$ is a SMS. Contradiction.
@@
(ii) If there exists a unique $c_{0}\in\mathcal{C}\setminus{\{w\}}$ s.t. $\delta(c_{0},{E}^{w\rightarrow})\leq 0$, taking ${E}^{w\rightarrow}$ and adding just enough element of ${E}^{\rightarrow c_{0}}$ so that $\delta(c_{0},\mathcal{X})=0$ gives a SMS (correctness of Algorithm 2). We just have to measure by how much we can ”deviate” from this SMS by swapping out elements of ${E}^{w\rightarrow}$ for elements of ${E}^{\rightarrow c_{0}}$ to keep $\delta(c_{0},\mathcal{X})=0$. Basically, by each swap out we reduce every score margin by at least 1. So at best, it is doable while the second smallest score margin remains positive i.e. $\min_{c\in\mathcal{C}\setminus{\{w,c_{0}\}}}\delta(c,{E}^{w\rightarrow})$ times. Note that this bound is not necessarily tight because we have to pick a specific candidate $c_{1}\neq c_{0}$ whose score margin would decrease by 2 instead of 1 when $(w,c_{1})$ gets swapped out.
@@
(iii) Here, $\forall c\neq w,\,\delta(c,{E}^{w\rightarrow})\geq 0$ and Algorithm 2 returns a SMS $\mathcal{X}\subseteq{E}^{w\rightarrow}$ s.t. $\max_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c_{0},\mathcal{X})\leq 1\lor\min_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c,\mathcal{X})\leq 0$. If $\min_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c,\mathcal{X})\leq 0$, we work our way with arguments similar to those introduced in (ii). If $\max_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c_{0},\mathcal{X})\leq 1$, $\sum_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c_{0},\mathcal{X})\leq m-1$. We recall that each element of ${E}^{w\rightarrow}$ contributes $m$ to $\sum_{c\neq w}\delta(c,\mathcal{X})$. Hence, only one swap out of an element of ${E}^{w\rightarrow}$ is possible before having a negative score margin for one candidate (decreasing the sum by $m$ and increasing it by $1$ with the swap in of an element not in ${E}^{w\rightarrow}$). ∎
@@
As we have shown that any SMS returned by Algorithm 2 is optimal regarding win count, it suffices to show that the maximum difference in terms of wins between the outcome of Algorithm 2 and any other SMS is smaller than $\frac{n}{2}$. The proof of this proposition uses the characterisation of SMSs for the Borda rule shown in Lemma 32. In case (i), the maximum difference is 0. In case (iii), it is $1$. In case (ii), it is at most $\min_{c\in\mathcal{C}\setminus{\{w,c_{0}\}}}\delta(c,{E}^{w\rightarrow})$. Let us show that this is less than $\frac{n}{2}$. Suppose that $\exists!c_{0}\in\mathcal{C}\setminus{\{w\}}$ s.t. $\delta(c_{0},{E}^{w\rightarrow})\leq 0$, let $\mathcal{X}$ be the SMS returned by Algorithm 2. We have that ${E}^{w\rightarrow}\subseteq\mathcal{X}$. Thus, by minimality of a MS we have $\forall c\in\mathcal{C}\setminus{\{w\}}$$\forall c^{\prime}\in\mathcal{C}\setminus{\{w,c_{0}\}}$, $\mu(c,c^{\prime})=0$. Hence, $\forall c\in\mathcal{C}\setminus{\{w,c_{0}\}}$, $\delta(c,\mathcal{X})=\sum_{i}\mu(w,i)+\sum_{i}\mu(i,c)-n(m-1)=\sum_{i}\mu(w,i)+\mu(w,c)-n(m-1)$ and $\delta(c_{0},{E}^{w\rightarrow})\leq 0$ can be rewritten as $\sum_{i\neq c_{0}}\mu(w,i)+2\mu(w,c_{0})\leq n(m-1)$. Additionally, $\sum_{i\neq c_{0}}\mu(w,i)\leq n(m-2)$. Hence,
@@
When $w$ is a Condorcet winner, we are in the second case of Proposition 15. Indeed, in that case, $\sigma_{w}=m-1$, $\min_{c\neq w}\mu(w,c)=1$ and we have m1>1(m1)1m-1>1(m-1)-1. Additionally, $\left\lceil 1(1-1/m\right\rceil=0$ and $\mathcal{X}=m-1$. The other case is clear. ∎
@@
Algorithm 3 is a simplified version of Algorithm 2, where we only consider the first branch of the if on line 4. We only reach the second branch when $w$ is a Condorcet winner and in that case $\{(w,c):c\in\mathcal{C}\}$ is a SMS. ∎
@@ 1 $\mathcal{X}\leftarrow{E}^{w\rightarrow}$
Appendix B Processes for Section 5
@@
$\operatorname{BO}$ aabbccdd3322443322221111 aa332244bb332211cc223311dd4422 $a$ is part of the Borda winners because $a$ wins at least 9=3+2+4 pairwise comparisons ($\mu_{\mathcal{X}}(a,b)=3$, $\mu_{\mathcal{X}}(a,b)=2$, $\mu_{\mathcal{X}}(a,b)=4$) $b$ wins at most 9=5*3-6 pairwise comparisons since it loses at least 6=3+2+1 ($\mu_{\mathcal{X}}(a,b)=3$, $\mu_{\mathcal{X}}(c,b)=2$, $\mu_{\mathcal{X}}(d,b)=1$) $c$ wins at most 9=5*3-6 pairwise comparisons since it loses at least 6=2+3+1 ($\mu_{\mathcal{X}}(a,c)=2$, $\mu_{\mathcal{X}}(b,c)=3$, $\mu_{\mathcal{X}}(b,c)=1$) $b$ wins at most 9=5*3-6 pairwise comparisons since it loses at least 6=4+2 ($\mu_{\mathcal{X}}(a,d)=4$, $\mu_{\mathcal{X}}(b,d)=2$)

### v3

@@ 11          $\mu_{\mathcal{X}}(c^{\prime},c)\leftarrow n-t$
Let $\mathcal{X}$ be an SMS for $w\in\operatorname{BO}(G)$. With Lemma 27, we show that if $\sigma_{w}\leq n(m-1)-\min_{c\neq w}\mu(w,c)$ then the lower bound of Lemma 24 is reached and for all SMSs $\mathcal{X}$ for $w\in\operatorname{BO}(G)$, $|\mathcal{X}|{=}(m-1)(n(m-1)-\sigma_{w})$.
@@
The lower bound directly result from the Theorem 15 as $n(m-1)-\lfloor n(1-1/m)\rfloor=n(m-1)-\left\lfloor n(m-1)/m\right\rfloor=\left\lceil n(m-1)-n(m-1)/m\right\rceil=\left\lceil\frac{n(m-1)^{2}}{m}\right\rceil$. For the upper bound, note that the Borda score of the winner $\sigma_{w}$ is greater than the average score of every candidate, i.e., $\sigma_{w}\geq\frac{1}{m}n\frac{m(m-1)}{2}$. In the worst case $\sigma_{w}=\left\lceil n\frac{m-1}{2}\right\rceil$ and $|\mathcal{X}|=(m-1)(n(m-1)-\left\lceil n\frac{m-1}{2}\right\rceil)=(m-1)\left\lfloor n\frac{m-1}{2}\right\rfloor$. ∎
@@
To compute a maxwin-SMS in $\mathcal{O}(m^{2}\log n)$, we propose the following Algorithm 2 which can naturally be derived from the analysis carried in the proof of Theorem 15 on the size of SMSs for the Borda rule. ∎
@@ 20    $\text{rmed}\leftarrow\text{rmed}+\text{iteRm}$
Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},E)$ , and a winning candidate $w\in\operatorname{BO}(G)$, for all SMSs $\mathcal{X}$ for $w\in\operatorname{BO}(G)$,
@@
    else if $\exists!c_{0}\in\mathcal{C}\setminus{\{w\}}$ s.t. $\delta(c_{0},{E}^{w\rightarrow})\leq 0$ then $|{E}^{w\rightarrow}\setminus{\mathcal{X}}|\leq\min_{c\in\mathcal{C}\setminus{\{w,c_{0}\}}}\delta(c,{E}^{w\rightarrow})$ and exists an SMS containing ${E}^{w\rightarrow}$
@@
    else for all SMSs $\mathcal{Y}$ for $w\in\operatorname{BO}(G)$, $|\operatorname{WC}(\mathcal{X})-\operatorname{WC}(\mathcal{Y})|\leq 1$.
@@
(i) We proceed by contradiction. Suppose, there is an SMS $\mathcal{X}$ of $w\in\operatorname{BO}(G)$ such that ${E}^{w\rightarrow}\not\subseteq\mathcal{X}$. Then, $\exists c_{2}\in\mathcal{C}\setminus{\{w\}}$ such that $(w,c_{2})\in E$ and $(w,c_{2})\not\in\mathcal{X}$. Additionally, we know that $\exists e_{0}\in{\mathcal{X}}^{\rightarrow c_{0}}$ such that $e_{0}\neq(w,c_{0})$ (all $c_{0}$’s defeats in $\mathcal{X}$ cannot come from $w$ since $\delta(c_{0},{E}^{w\rightarrow})<0$). Similarly, $\exists e_{1}\in{\mathcal{X}}^{\rightarrow c_{1}}$, $e_{1}\neq(w,c_{1})$. Let $\mathcal{X}^{\prime}=\mathcal{X}\setminus{\{e_{0},e_{1}\}}\cup\{(w,c_{2})\}$. $\forall c\in\mathcal{C}\setminus{\{w,c_{0},c_{1},c_{2}\}},\,\delta(c,\mathcal{X}^{\prime})=\delta(c,\mathcal{X})+1\geq 0$. $\delta(c_{2},\mathcal{X}^{\prime})=\delta(c_{2},\mathcal{X})+1\geq 0$. $\forall c\in\{c_{0},c_{1}\}\setminus{\{c_{2}\}}$, $\delta(c,\mathcal{X}^{\prime})=\delta(c,\mathcal{X})-1+1\geq 0$. Hence $\mathcal{X}^{\prime}$ is a weak MS and it contains at least one MS with a smaller cardinality. However, $|\mathcal{X}^{\prime}|<|\mathcal{X}|$ and $\mathcal{X}$ is an SMS. Contradiction.
@@
(ii) If there exists a unique $c_{0}\in\mathcal{C}\setminus{\{w\}}$ s.t. $\delta(c_{0},{E}^{w\rightarrow})\leq 0$, taking ${E}^{w\rightarrow}$ and adding just enough element of ${E}^{\rightarrow c_{0}}$ so that $\delta(c_{0},\mathcal{X})=0$ gives an SMS (correctness of Algorithm 2). We just have to measure by how much we can ”deviate” from this SMS by swapping out elements of ${E}^{w\rightarrow}$ for elements of ${E}^{\rightarrow c_{0}}$ to keep $\delta(c_{0},\mathcal{X})=0$. Basically, by each swap out we reduce every score margin by at least 1. So at best, it is doable while the second smallest score margin remains positive i.e. $\min_{c\in\mathcal{C}\setminus{\{w,c_{0}\}}}\delta(c,{E}^{w\rightarrow})$ times. Note that this bound is not necessarily tight because we have to pick a specific candidate $c_{1}\neq c_{0}$ whose score margin would decrease by 2 instead of 1 when $(w,c_{1})$ gets swapped out.
@@
(iii) Here, $\forall c\neq w,\,\delta(c,{E}^{w\rightarrow})\geq 0$ and Algorithm 2 returns an SMS $\mathcal{X}\subseteq{E}^{w\rightarrow}$ s.t. $\max_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c_{0},\mathcal{X})\leq 1\lor\min_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c,\mathcal{X})\leq 0$. If $\min_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c,\mathcal{X})\leq 0$, we work our way with arguments similar to those introduced in (ii). If $\max_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c_{0},\mathcal{X})\leq 1$, $\sum_{c\in\mathcal{C}\setminus{\{w\}}}\delta(c_{0},\mathcal{X})\leq m-1$. We recall that each element of ${E}^{w\rightarrow}$ contributes $m$ to $\sum_{c\neq w}\delta(c,\mathcal{X})$. Hence, only one swap out of an element of ${E}^{w\rightarrow}$ is possible before having a negative score margin for one candidate (decreasing the sum by $m$ and increasing it by $1$ with the swap in of an element not in ${E}^{w\rightarrow}$). ∎
@@
As we have shown that any SMS returned by Algorithm 2 is optimal regarding win count, it suffices to show that the maximum difference in terms of wins between the outcome of Algorithm 2 and any other SMS is smaller than $\frac{n}{2}$. The proof of this proposition uses the characterisation of SMSs for the Borda rule shown in Lemma 32. In case (i), the maximum difference is 0. In case (iii), it is $1$. In case (ii), it is at most $\min_{c\in\mathcal{C}\setminus{\{w,c_{0}\}}}\delta(c,{E}^{w\rightarrow})$. Let us show that this is less than $\frac{n}{2}$. Suppose that $\exists!c_{0}\in\mathcal{C}\setminus{\{w\}}$ s.t. $\delta(c_{0},{E}^{w\rightarrow})\leq 0$, let $\mathcal{X}$ be the SMS returned by Algorithm 2. We have that ${E}^{w\rightarrow}\subseteq\mathcal{X}$. Thus, by minimality of an MS we have $\forall c\in\mathcal{C}\setminus{\{w\}}$$\forall c^{\prime}\in\mathcal{C}\setminus{\{w,c_{0}\}}$, $\mu(c,c^{\prime})=0$. Hence, $\forall c\in\mathcal{C}\setminus{\{w,c_{0}\}}$, $\delta(c,\mathcal{X})=\sum_{i}\mu(w,i)+\sum_{i}\mu(i,c)-n(m-1)=\sum_{i}\mu(w,i)+\mu(w,c)-n(m-1)$ and $\delta(c_{0},{E}^{w\rightarrow})\leq 0$ can be rewritten as $\sum_{i\neq c_{0}}\mu(w,i)+2\mu(w,c_{0})\leq n(m-1)$. Additionally, $\sum_{i\neq c_{0}}\mu(w,i)\leq n(m-2)$. Hence,
@@
When $w$ is a Condorcet winner, we are in the second case of Theorem 15. Indeed, in that case, $\sigma_{w}=m-1$, $\min_{c\neq w}\mu(w,c)=1$ and we have m1>1(m1)1m-1>1(m-1)-1. Additionally, $\left\lceil 1(1-1/m\right\rceil=0$ and $\mathcal{X}=m-1$. The other case is clear. ∎
@@
Algorithm 3 is a simplified version of Algorithm 2, where we only consider the first branch of the if on line 4. We only reach the second branch when $w$ is a Condorcet winner and in that case $\{(w,c):c\in\mathcal{C}\}$ is an SMS. ∎
@@ 1 $\mathcal{X}\leftarrow{E}^{w\rightarrow}$
Appendix B Additional Examples for Section 5
@@
$\operatorname{BO}$ aabbccdd3322443322221111 aa332244bb332211cc223311dd4422 $a$ is part of the Borda winners because $a$ wins at least 9=3+2+4 pairwise comparisons ($\mu_{\mathcal{X}}(a,b)=3$, $\mu_{\mathcal{X}}(a,c)=2$, $\mu_{\mathcal{X}}(a,d)=4$) $b$ wins at most 9=5*3-6 pairwise comparisons since it loses at least 6=3+2+1 ($\mu_{\mathcal{X}}(a,b)=3$, $\mu_{\mathcal{X}}(c,b)=2$, $\mu_{\mathcal{X}}(d,b)=1$) $c$ wins at most 9=5*3-6 pairwise comparisons since it loses at least 6=2+3+1 ($\mu_{\mathcal{X}}(a,c)=2$, $\mu_{\mathcal{X}}(b,c)=3$, $\mu_{\mathcal{X}}(d,c)=1$) $d$ wins at most 9=5*3-6 pairwise comparisons since it loses at least 6=4+2 ($\mu_{\mathcal{X}}(a,d)=4$, $\mu_{\mathcal{X}}(b,d)=2$)